# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

### v1

@@ Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. We demonstrate that this interplay gives rise to a new class of mixed-state phase transitions, using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional dimerized Heisenberg model. Furthermore, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional Hamiltonians. Our results highlight MDITE as a powerful paradigm for investigating non-unitary dynamics and the fundamental role of decoherence in many-body quantum systems.
@@ I Introduction
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [Garratt2023nonlocal-measure, Nahum2020majoranadecoherence, Lee2023decoh_criticality, Amin2024symmetry-mixed-ent, Passarelli2024mixed-mite, Su2024higherform-symmetry, Tantivasadakarn2024longrange-ent-measure, Fanruihua2024diagnostics, wangzhu2025decoherence]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
@@
Specifically, our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a many-body Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the stationary state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian. As we will demonstrate in Sec. IV and V, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state phase transitions, which can be characterized using conventional linear order parameters—such as spin magnetization—without relying on information-theoretic proxies.
@@
This paper is organized as follows. In Sec. II, we formally define the MDITE protocol and highlight its key features. In Sec. III, we provide a schematic representation of the MDITE process, which not only clarify the setup but also facilitate numerical studies including quantum Monte Carlo (QMC) methods. The one- and two-dimensional Hamiltonians studied in this work, along with the corresponding numerical results for mixed-state phases and criticalities, are introduced and analyzed in Secs. IV and V. Finally, in Sec. VI, we summarize the work and discuss several promising directions for future research.
@@ IV.5 Analysis of the criticalities
Building upon the consistent critical exponent ratio $\beta/\nu\approx 0.40$ observed across all dynamical transitions in our simulations, and supported by the unified dynamical mechanism described above, we argue that the mixed-state phase transitions related to the 1D TFIM fall within a single universality class when $p>0$. This universality indicates the existence of a continuous critical surface in the extended parameter space $(\tau,h,p)$ separating the paramagnetic and ferromagnetic phases.
@@
In addition, the emergence of long-range correlations leading to the Ising-like phase with increasing measurement rate can also be understood qualitatively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites during the QMC updates. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. C.
@@ V.3 Mixed-state phase transitions and analysis
To interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the $SU(2)$ symmetry, favoring the $z$-axis. Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. When $g>g_{c,\mathrm{GS}}$, and the imaginary-time steps are short ($\tau g\ll 1$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gg 1$, the off-diagonal coherence builds up rapidly, and the final stationary state remains disordered even at $p=1$. Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM effectively generalizes the physics of the 1D TFIM to two dimensions.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range correlations induced by measurements can naturally be understood qualitatively through the cluster formation mechanism, as discussed in Appendix. C. We emphasize that it would be interesting to explore other types of decoherence channels that preserve the $U(1)$ or $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC stimulations would be necessary, potentially revealing qualitatively different types of phase transitions. Physically, the enhancement of long-range order induced by measurement can be interpreted as a suppression of quantum fluctuations—i.e., the off-diagonal elements of the density matrix—caused by the diagonal measurement employed in our setup.
@@ VI Conclusion and discussions
Our work opens several promising avenues for future research. One direction is to further characterize the universality classes associated with the criticalities observed in our simulations. Moreover, our QMC approach allow systematic investigations of dynamical properties and nonequilibrium phenomena of the MDITE process. Another is to extend the framework to more complex Hamiltonians and general decoherence channels beyond projective measurements, thereby accessing a broader landscape of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [lu2023charac, lu2023longrange, Fanruihua2024diagnostics, Lu2024disentangling, lessa2025higherformanomaly], as well as the strong-to-weak spontaneous symmetry-breaking [Lee2023decoh_criticality, Lessa2025swssb, sala2024swssb, gu2024swssb, huang2025swssb, zack2025renyi1corr, guo2025swssb, guoyc2025swssb, guoyc2025ldpo, zhang2025swssb] can be explored within the MDITE framework.
@@ Acknowledgements
We thank G.-Y. Zhu, B.-B. Miao and Z. Bi for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
@@ Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
Appendix C Cluster formations and long-range correlations
In this section, from the perspective of QMC simulations, we provide an intuitive understanding of how measurements induce long-range correlations in the system, leading to the emergence of a mixed and ordered phase.
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, larger clusters automatically formed during the update process correspond to stronger correlations between different sites. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. This extensive connectivity generates strong correlations throughout the system, corresponding to the ferromagnetic phase with long-range order.
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, even starting with the paramagnetic ground state of the TFIM, measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This is the reason why measurements can drive the system into a mixed-ordered phase with long-range correlations, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
For general Hamiltonians, the specific form of the nonlocal update scheme may differ from the cluster update employed in the TFIM and CDHM. Depending on the model, the corresponding update mechanism could also reveal new insights into the intrinsic properties of the states in the MDITE process.

### v2

@@ Abstract
Mixed-state phase transitions have recently attracted growing attention as a new frontier in nonequilibrium quantum matter and quantum information. In this work, we introduce the measurement-dressed imaginary-time evolution (MDITE) as a novel framework to explore mixed-state quantum phases and decoherence-driven criticality. In this setup, alternating imaginary-time evolution and projective measurements generate a competition between coherence-restoring dynamics and decoherence-inducing events. While reminiscent of monitored unitary circuits, MDITE fundamentally differs in that the physics is encoded in decoherent mixed states rather than in quantum trajectories. We demonstrate that this interplay gives rise to a novel class of mixed-state phase transitions, using numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model. Notably, the observed transitions do not fall into any previously established universality classes. Furthermore, we provide a diagrammatic representation of the evolving state, which naturally enables efficient studies of MDITE with quantum Monte Carlo and other many-body numerical methods, thereby extending investigations of mixed-state phase transitions to large-scale and higher-dimensional systems. In addition, the representation provides a natural interpretation of the phase transitions in terms of cluster formation within the simulations. Our results highlight MDITE as a powerful paradigm for investigating non-unitary dynamics and the fundamental role of decoherence in many-body quantum systems.
@@ I Introduction
A complementary, yet fundamentally distinct, perspective arises by considering the measurement-averaged mixed state, which directly captures the universal characterstics of decoherence [Garratt2023nonlocal-measure, Nahum2020majoranadecoherence, Lee2023decoh_criticality, Amin2024symmetry-mixed-ent, Passarelli2024mixed-mite, Su2024higherform-symmetry, Tantivasadakarn2024longrange-ent-measure, Fanruihua2024diagnostics, wangzhu2025decoherence, guoyc2025lindladian, guoyc2025swssb, guoyc2025ldpo]. In contrast to conventional MIPTs which are fundamentally rooted in quantum trajectories of monitored unitary circuits, we instead consider a mixed-state perspective based on the interplay between imaginary-time evolution (ITE) and projective measurements—two intrinsically non-unitary processes governing the dynamics. Building on this idea,, we introduce the measurement-dressed imaginary-time evolution (MDITE) (Sec. II).
@@
Specifically, our MDITE setup begins with a maximally mixed state and alternates between ITE controlled by a many-body Hamiltonian and random projective measurements. In the absence of measurement, ITE acts as a deterministic mechanism, steering the system toward the ground state of the Hamiltonian. Over extended times, this evolution suppresses classical uncertainty while reinforcing quantum coherence. In contrast, nonzero projective measurements tend to amplify classical uncertainty and induce decoherence, counteracting the ordering effects of ITE. This competition suggests the existence of mixed-state phase transitions with respect to the stationary state of the process driven by the strength of decoherence, evolution time, and parameters in the Hamiltonian. As we will demonstrate in Sec. IV and V, through large-scale numerical simulations of the one-dimensional transverse-field Ising model and the two-dimensional columnar dimerized Heisenberg model, we reveal the existence of such mixed-state phase transitions, which can be characterized using conventional linear order parameters—such as spin magnetization—without relying on information-theoretic proxies. Notably, our detailed analysis of the critical behavior shows that, to the best of our knowledge, these transitions do not fall into any previously established universality classes.
@@
This paper is organized as follows. In Sec. II, we formally define the MDITE protocol and highlight its key features. In Sec. III, we provide a schematic representation of the MDITE process, which not only clarify the setup but also facilitate numerical studies including quantum Monte Carlo (QMC) methods. Within this representation, the observed mixed-state phase transitions admit a natural interpretation in the picture of cluster formation in the QMC updates. The one- and two-dimensional Hamiltonians studied in this work, along with the corresponding numerical results for mixed-state phases and criticalities, are introduced and analyzed in Secs. IV and V. Finally, in Sec. VI, we summarize the work and discuss several promising directions for future research.
@@ IV.5 Analysis of the criticalities
Building upon the consistent critical exponent ratio $\beta/\nu\approx 0.40$ observed across all dynamical transitions in our simulations, and supported by the unified dynamical mechanism described above, we argue that the mixed-state phase transitions related to the 1D TFIM fall within a single universality class when $p>0$. This universality indicates the existence of a continuous critical surface in the extended parameter space $(\tau,h,p)$ separating the paramagnetic and ferromagnetic phases. According to the finite-size scaling hypothesis and renormalization-group theory [fisher1972scaling, sandvik2010computational, Wilson1971Renormalization, Cardy1996], the magnetization near a critical point obeys the scaling form $\langle m\rangle\sim L^{-\beta/\nu}f(gL^{1/\nu})$, where $g$ is the reduced parameter around the critical point and $f$ is a dimensionless function. In the nonequilibrium universality class considered here, the ratio $\beta/\nu$ remains universal, whereas the individual critical exponents $\beta$ and $\nu$ can depend on the path along which the critical point is approached. Consequently, different parameter trajectories may exhibit distinct values of $\beta$ and $\nu$, while their ratio $\beta/\nu$, which governs the leading finite-size scaling of the magnetization, remains the same [wang2025sudden].
@@
In addition, the emergence of long-range order with an increasing measurement rate can also be understood qualitatively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites during the QMC updates. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. C.
@@ V.3 Mixed-state phase transitions and analysis
To interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the $SU(2)$ symmetry to a residual $U(1)\rtimes\mathbb{Z}_{2}$, favoring the $z$-axis. According to the Mermin–Wagner theorem [Mermin1966, Hohenberg1967, Sondhi1997Continuous], a continuous $U(1)$ symmetry cannot undergo spontaneous symmetry breaking in two dimensions for any finite imaginary-time extent. Consequently, the transition observed here can only break the discrete symmetry $\mathbb{Z}_{2}$ induced by the projective measurements. Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. In Appendix. D, we present the results demonstrating the absence of off-diagonal long-range order in this model.
When $g>g_{c,\mathrm{GS}}$, and the imaginary-time steps are short ($\tau g\ll 1$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gg 1$, the off-diagonal coherence builds up rapidly, and the final stationary state remains disordered even at $p=1$. Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM effectively generalizes the physics of the 1D TFIM to two dimensions.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range orders induced by measurements can naturally be understood qualitatively through the cluster formation mechanism, as discussed in Appendix. C. We emphasize that it would be interesting to explore other types of decoherence channels that preserve other symmetries, such as the $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC stimulations would be necessary, potentially revealing qualitatively different types of phase transitions. Physically, the enhancement of long-range order induced by measurement can be interpreted as a suppression of quantum fluctuations—i.e., the off-diagonal elements of the density matrix—caused by the diagonal measurement employed in our setup.
@@ VI Conclusion and discussions
Our work opens several promising avenues for future research. One direction is to further characterize the universality classes associated with the criticalities observed in our simulations. Moreover, our QMC approach allow systematic investigations of dynamical properties and nonequilibrium phenomena of the MDITE process. Another is to extend the framework to more complex Hamiltonians and general decoherence channels beyond projective measurements, thereby accessing a broader landscape of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [lu2023charac, lu2023longrange, Fanruihua2024diagnostics, Lu2024disentangling, sang2025markov, lessa2025higherformanomaly], as well as the strong-to-weak spontaneous symmetry-breaking [Lee2023decoh_criticality, Lessa2025swssb, sala2024swssb, gu2024swssb, huang2025swssb, zack2025renyi1corr, guo2025swssb, guoyc2025swssb, guoyc2025ldpo, zhang2025swssb, zpf2025swssb, Liu2025swssb] can be explored within the MDITE framework.
@@ Acknowledgements
We thank G.-Y. Zhu, B. Miao and Z. Bi for helpful discussions. Zenan Liu thanks the China Postdoctoral Science Foundation under Grants No.2024M762935 and NSFC Special Fund for Theoretical Physics under Grants No.12447119. Zhe Wang thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. This project is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Data availability. Data are available on Zenodo [zenodo].
@@ Appendix B Quantum Monte Carlo algorithm for simulating the generalized partition function $\mathcal{Q}_{n_{d}}$
Appendix C Cluster formations and long-range orders
In this section, we provide an intuitive, QMC-based picture for how local projective measurements enhance collective behavior in the system, ultimately generating a mixed ordered phase.
As introduced in Appendix. A, the nonlocal cluster update is a crucial step to ensure the ergodicity of the QMC algorithm for the TFIM. Therefore, the size of the automatically formed clusters reflects the degree of collective behavior among spins: large clusters emerge when many spins fluctuate coherently. For example, in the 1D TFIM, when the transverse field is very strong $(h\to\infty)$, spins are dominated by the external field, and cluster updates produce only small, local clusters. Conversely, when the field is very weak, $(h\to 0)$, the Ising interactions dominate, connecting all sites into a single large clusters. In this regime, cluster construction naturally produces system-spanning clusters, reflecting the collective alignment of spins in the ferromagnetic phase.
According to the merge-split process introduced in Appendix. B, the presence of measurements allows update lines to connect across replicas, effectively merging clusters from different replicas. Therefore, starting with the paramagnetic ground state of the TFIM, even local measurements can induce the formation of larger clusters during the QMC updates, which is similar to increasing the Ising coupling strength for the ground state of the TFIM. This is the reason why measurements can drive the system into a mixed-ordered phase with long-range order, from the perspective of cluster formations in QMC simulations. Similar reasoning also applies to the CDHM, where a cluster update is also used.
We emphasize that for general Hamiltonians, the specific form of the nonlocal update scheme may differ from the cluster update employed in the TFIM and CDHM. Depending on the model, the corresponding update mechanism could also reveal new insights into the intrinsic properties of the states in the MDITE process.
Appendix D Results on the absence of off-diagonal long-range order in the 2D CDHM
[…]
Figure 13: Finite size extrapolation of off-diagonal spin-spin correlation $C_{XY}(L/2,L/2)$ at different measurement probability $p$ with $\tau=16$ (a) and $\tau=100$ (b).
In this section, we present additional numerical results demonstrating the absence of off-diagonal long-range order in the 2D CDHM under the MDITE protocol, as discussed in Sec. V. Combining the algorithm in Sec. C and the QMC tomography method [mao2025sampling, hari2025magic, chincholi2025detecting], we compute the correlation function
$C_{XY}(\textbf{r})=\langle X_{\textbf{r}}X_{0}+Y_{\textbf{r}}Y_{0}\rangle.$ (19)
To extrapolate to the limit $|\textbf{r}|\to\infty$, we evaluate $C_{XY}(\textbf{r})$ at the maximum separation $\textbf{r}=(L/2,L/2)$ for each system size $L\times L$. We perform a finite-size extrapolation of $C_{XY}(L/2,L/2)$ as a function of $1/L$. As shown in Fig. 13, for sufficiently long imaginary time $\tau$, $C_{XY}(L/2,L/2)$ extrapolates to zero in the thermodynamic limit ($L\to\infty$), indicating the absence of off-diagonal long-range order in the mixed states generated by the MDITE protocol.